x^2-8x+15=0
how do you solve this problem, and how do you use the discriminant?

Answer by checkley77(12844)

x^2-8x+15=0
(x-5)(x-3)=0
x-5=0
x=5 ans.
x-3=0
x=3 ans.
The discriminate is b^2-4ac.
-8^2-4*1*15=-64-60=4
Because the answer is a perect square then the answers are 2 real rational numbers.